OpenStudy (anonymous):

which of the following is the solution to the differential equation dy/dx=y^2 where y(-1)=1? a)y=1/x for x cannot be = 0 b)y=-1/x for x<0 c)y=-1/x for x>0 d)y=1/x for x>0 e)y=1/x for x<0

OpenStudy (anonymous):

\[\frac{dy}{dx}=y^2\\ \frac{1}{y^2}~dy=dx\] Integrating both sides yields \[-\frac{1}{y}=x+C~~\iff~~y=-\frac{1}{x+C}\] Now, given that \(y(-1)=1\), you have \[1=-\frac{1}{-1+C}~~\Rightarrow~~C=0\]

OpenStudy (anonymous):

what would the answer be

OpenStudy (anonymous):

I'd go for b

OpenStudy (anonymous):

thank you:)